OpenStudy (tortillabluechip):

Which of the following options is an equivalent function to f(x) = 3(2)^3x a. f(x) = 24^x b. f(x) = 27(8)^x c. f(x) = 3(8)^x d. f(x) = 3(8x)

OpenStudy (mathmale):

Word to the wise: Please express exponentiation properly: Correct: f(x) = 3(2)^(3x) Vague: f(x) = 3(2)^3x Note: \[a ^{bc}=(a^b)^c\] Can you use this fact to answer the question at hand?